OpenStudy (anonymous):

Betty bought 4 beads for 22 cents. The red beads were 10 cents each and the blue beads were 1 cent each. Find the number of red and blue beads.

OpenStudy (mathstudent55):

Do you need to solve with an equation, or a system of equations, or is any method ok?

OpenStudy (anonymous):

Any method is ok

OpenStudy (mathstudent55):

Since the purcahse is only 4 items, we can list all possibilities: Red beads Cost Blue beads Cost Total Cost Number bought 0 0 4 4 4 Number bought 1 10 3 3 13 Number bought 2 20 2 2 22 Number bought 3 30 1 1 31 Number bought 4 40 0 0 40

OpenStudy (anonymous):

2 red & 2 blue X+y = 4 10x + 1y =22 10x + 4-x=22 9x=18 X=2.

OpenStudy (anonymous):

two red and two blue

OpenStudy (mathstudent55):

correct